OpenStudy (anonymous):

Find the slope of the line -9x-5y=3

OpenStudy (anonymous):

put in slope intercept form y=mx+b and find m(coefficient of x-number in from of x))

OpenStudy (anonymous):

5y=-9x-3 /5 /5 /5 y=-9/5x-3/5 m=-9/5

OpenStudy (anonymous):

you got it